A vertical spring with a spring constant of 460 N/m is mounted on the floor. From directly above the spring, which is unstrained, a 0.30 kg block is dropped from rest. It collides with and sticks to the spring, which is compressed by 2.3 cm in bringing the block to a momentary halt. Assuming air resistance is negligible, from what height (in cm) above the compressed spring was the block dropped?
Answer #1

Assuming air resistance is negligible, from what height (in cm) above the compressed spring was the block dropped?

Using conservation of energy, we have

P.Egravity = P.E​​​​spring

m g h = (1/2) k x​​​​​​2

(0.30 kg) (9.8 m/s​​​​​​2​​​) h = (0.5) (460 N/m) (0.023 m)2

(2.94 N) h = (0.12167 N.m)

h = [(0.12167 N.m) / (2.94 N)]

h = 0.0413 m

Converting m into cm :

h = 4.13 cm